OpenStudy (anonymous):

-2(45-5)-10=-3(6s+4)-3

OpenStudy (anonymous):

options?

OpenStudy (anonymous):

solve the equation

OpenStudy (anonymous):

s = 25/6 -90 = -18 s-15

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

17(t-5)+4t=7(3t+3)-12 solve the equation